Three identical units of merchandise were purchased during July, as follows: Date Product T Units Cost

July 3 Purchase 1 $20.00

10 Purchase 1 23.00

24 Purchase 1 26.00

Total 3 $69.00 Average cost per unit $23.00 Assume one unit sells on July 28 for $34.00.

Determine the gross profit, cost of merchandise sold, and ending inventory on July 31 using the (a) first-in, first-out, (b) last-in, first-out, and (c) average cost flow methods.
Answer #1
FIFO LIFO Average cost
Ending inventory 23+26 = 49 20+23 = 43 23*2 = 46
Cost of goods sold 20 26 23
Gross profit (Sales-Cost of goods sold) 34-20 = 14 34-26 = 8 34-23 = 11